OpenStudy (anonymous):

multiply and simplify in lowesst terms then put the answer in factor form Please help the problem I will do on the equation

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ x ^{2}+11 }*\frac{ x+8 }{ x}\]

OpenStudy (tyteen4a03):

Cross out the pair of x^2 and x so it becomes \(\frac{x(x+8)}{x^2+11}\). (alternatively turn x(x+8) into x^2 + 8.

OpenStudy (tyteen4a03):

Uh, I mean x^2 + 8x

OpenStudy (anonymous):

ok got that now what do I do

OpenStudy (tyteen4a03):

That would be the final answer, given that you're only given an expression.

OpenStudy (anonymous):

what about the 11 at the bottom that soes not make sunce to me

OpenStudy (anonymous):

HELP

OpenStudy (tyteen4a03):

The x^2 + 11 at the bottom was never modified (because you can't factor it without bringing out complex numbers)

OpenStudy (anonymous):

So you are saying the factor form is x^2+8x?

OpenStudy (anonymous):

tyteen4a03?

OpenStudy (tyteen4a03):

One of the answers would be x^2 + 8x, but it isn't the factored form (IMO the factor form is meaningless in this question) The factor form of x^2 + 8x would be x(x+8).

OpenStudy (tyteen4a03):

As for the most simplified form of the expression given, it's \(\frac{x^2+8x}{x^2+11}\)

OpenStudy (anonymous):

k thanks